| 12/22/2004 | |||||||
| 1:08 | data flow didn't connect yet | JH | |||||
| ESC said the satellite is on the horizen | |||||||
| 1:34 | received the data again | SO | |||||
| For some problems on SW, we restart the mornitoring program. | |||||||
| 1:51 | we tried to send "NOOP" command, but it didn't go through. | ||||||
| 1:55 | SCI_SW_OPMODE 0 | SO | |||||
| SCI_SNIO_CONTROL 21,0 | |||||||
| Execute script daq_restart.stol | |||||||
| 2:03 | SCI_SW_OPMODE 0 | SO | |||||
| SCI_SNIO_CONTROL 21,0 | |||||||
| Execute script daq_restart.stol | |||||||
| 2:13 | SCI_SW_OPMODE 0 | SO | |||||
| SCI_SNIO_CONTROL 21,0 | |||||||
| Execute script daq_restart.stol | |||||||
| 2:13 | sent command "SCI_SW_HV_LEVEL 0x14, 1000" | JH | |||||
| but "ERROR: HVC: :SetTargetLevel: invalid box 20" in GUI log | |||||||
| 2:18 | Network Rate: 1.3kbps, so data flow was very slow | JH | |||||
| 2:24 | SCI_SW_OPMODE 0 | SO | |||||
| SCI_SNIO_CONTROL 21,0 | |||||||
| Execute script daq_restart.stol | |||||||
| 2:33 | SCI_SW_OPMODE 0 | SO | |||||
| SCI_SNIO_CONTROL 21,0 | |||||||
| Execute script daq_restart.stol | |||||||
| sent stol file again | |||||||
| 2:45 | SCI_SW_OPMODE 0 | JH | |||||
| SCI_SNIO_CONTROL 21,0 | |||||||
| Execute script daq_restart.stol | |||||||
| 2:54 | clicked the "standby mode" button in GUI mode to reset the data receiving program, and then cleared the buffer. | SO | |||||
| SCI_SW_OPMODE 0 | |||||||
| SCI_SNIO_CONTROL 21,0 | |||||||
| Execute script daq_restart.stol | |||||||
| but the error had not been solved. | |||||||
| 3:02 | Z-HI trigger rate was too low from 20041222-011515.dat | JH | |||||
| continued to calibration trigger not Z-HI trigger | |||||||
| called Sonny and we sent some commands | |||||||
| stanby mode in GUI of cdps2 and sent above 3 commands | |||||||
| 3:21 | SCI_TCD_RESET | JH | |||||
| 3:29 | run 10000 evts of Master Trigger, CAL, HDS and SCD in GUI of cdps1 | JH | Data disconnected as of 3:30 | ||||
| 4:10 | housekeeping monitoring didn't update yet | JH | |||||
| ESC said they also didn't get any data from CDM | |||||||
| 7:17 | start getting data | HQ | |||||
| 8:05 | Command window comes back | ||||||
| I just tried previous command again | YS | ||||||
| SCI_SW_OPMODE 0 | |||||||
| SCI_SNIO_CONTROL 21,0 | |||||||
| Execute script daq_restart.stol | |||||||
| SCI_TCD_RESET | YS | ||||||
| Then, Z-HI trigger rate became normal about 20Hz | YS | ||||||
| Trigger rate looks ok, but still there are a lot of Warning messages for TCD fail to connect | YS | ||||||
| 8:36 | asked ESC for TCD power off | JH | |||||
| asked ESC for TCD power on again | |||||||
| 8:39 | sent command "CMD SCI_SW_DAQ_RESET" | JH | |||||
| but error massage came up, so tried again | |||||||
| 8:43 | SCI_SPARE1, 0 & scd_ch_mask_hot3.stol | JH | |||||
| cal_trig_threshold changed from 100 to 85 | |||||||
| arc_start.stol & daq_start.stol | |||||||
| 9:13 | but they didn't go through completely, so we tried again | JH | |||||
| 9:17 | TCD hv off & power off | ||||||
| 9:22 | SCD ch#2880 threshold level changed from 40000 to 70000 | JH | |||||
| 9:33 | asked ESC for TCD power off | ||||||
| 9:35 | SCD ch#2880 threshold level changed from 40000 to 70000 | JH | |||||
| but STV value in spi file became 4464->why? | |||||||
| 9:45 | Simon want to turn off TCD for a while to cool down | YS | |||||
| we asked ESC to turn off TCD Power | |||||||
| Send commands to disable TCD | |||||||
| SCI_SW_OPMODE 0 | |||||||
| SCI_SW_SELECT_EVENT 1, 0xBA | |||||||
| SCI_SW_DAQ_RESET | |||||||
| send "SCI_SPARE1 0" and "scd_ch_mask_hot3.stol" | |||||||
| DAQ_START.stol | |||||||
| 9:57 | ESC power on TCD | ||||||
| SCI_SW_SELECT_EVENT 1, 0xBE | |||||||
| SCI_SW_DAQ_RESET | |||||||
| send "SCI_SPARE1 0" and "scd_ch_mask_hot3.stol" | |||||||
| arc_start.stol | |||||||
| DAQ_START.stol | |||||||
| SCI_TCD_RESET | JH | ||||||
| 10:00 | TCD hv on: SCI_TCD_HV | ||||||
| 11:06 | Stand by button | YS | |||||
| DAQ reset button (check message: shouldn't have connection failure error messages, should have "Ready to Go" ) | |||||||
| ERROR | |||||||
| POWER_ON_TCD.stol (reset TCD thresholds) | |||||||
| HV_ON_TCD.stol | |||||||
| send "SCI_SPARE1 0" and "scd_ch_mask_hot3.stol" | |||||||
| DAQ_START.stol (confirm 0x00BE) | |||||||
| Remember trigger threshold resets to default value whenever you send "SCI_SPARE1 0" | |||||||
| hprint -tcd hskfile to see TCD hsk setting | |||||||
| which hprint | |||||||
| 12:00 noon | stop taking data -stand by mode | ||||||
| 12:04 | arc_stop.stol | ||||||
| power_off_cpu.stol (confirm SW_shutdown? 3 line messages) | |||||||
| Wait 3 min (HSK to reach 120 sec) | |||||||
| 12:07 | ESC to turn off SFC (SFC current reading should be 0) | ||||||
| ESC to turn on SFC | |||||||
| check TCD connections (check error messages for TCD connection before Ready to Go message) | |||||||
| 12:14 | send "SCI_SPARE1 0" and "scd_ch_mask_hot3.stol" | ||||||
| arc_start.stol | |||||||
| 12:15 | set CAL thrigger threshold to DAC 85 | ||||||
| 12:18 | DAQ_START.stol | ||||||
| powering off command box causes resetting cal boxes to lose HV | |||||||
| Must ramp down HV & bias before turning off SFC | |||||||
| Consider having ESC reset SFC before taking SFC power procedure | |||||||
| 12:51 | Power on SCD by "power_on_scd.stol" | ||||||
| Bias on SCD by "hv_on_scd.stol" | |||||||
| 12:54 | Turn on Bias of Calorimeter and Hodoscope | ||||||
| Send HV for Cal 2000 (6kV) | |||||||
| Send Cal threshold DAC value to 100 (during ramping) | YS | ||||||
| 13:05 | Send HV for HDS (3000) | YS | |||||
| 13:14 | Simon Set TRD HV1=1.8kV to track Z=1 | YS | |||||
| 13:20 | Calorimeter rampimg is done. | YS | |||||
| 13:39 | Hodoscope rampin is done. (GMT 18:39) | ||||||
| Data taken during ramping up is not reliable. | |||||||
| Data file from now should be good. | |||||||
| 1:46 | Send cal threshold of 82 | ||||||
| See operation manual for the Emergency TRD shut down procedure that Simon provided today | |||||||
| If TRD should be off/on, set TRD HV as following | |||||||
| HV 1 (TRD1) 1870V | |||||||
| HV 5 (TRD2) 1900 V | |||||||
| HV 9 (Cherenkov) 1225 V | |||||||
| Check HSK monitoring values should be as following | |||||||
| HV1 = 1.77 KV | |||||||
| HV2 = 1.82 KV | |||||||
| HV3 = 1.18 KV | |||||||
| If TCD should be off/on check if level 3 trigger is 300 mV. This is different from the default 350 mV. | |||||||
| 15:30 | Send cal threshold of 75 | HQ | |||||
| Send cal threshold of 72 | |||||||
| Send cal threshold of 70 | |||||||
| 15:45 | HOD MB3 HV 7.99 (even set the DAC value 3500) | HQ | |||||
| 17:04 | SCI_SW_SCD_ENABLE_AUSTV 0xFFF, 0, 3 | ||||||
| SCI_SW_CAL_PED_UPDATE_RUN 100 | |||||||
| >>> SCD sparsification didn't change. | |||||||
| 17:10 | To ramping up cal MB5, send "SCI_SW_HV_LEVEL 0x04, 250 | OG | |||||
| To ramping up H3, send "SCI_SW_HV_LEVEL 0x0F, 3750 | |||||||
| To ramping up cal MB5, send "SCI_SW_HV_LEVEL 0x04, 500 | |||||||
| To ramping up cal MB5, send "SCI_SW_HV_LEVEL 0x04, 2000 | |||||||
| To ramping up H8, send "SCI_SW_HV_LEVEL 0x14, 2000 | |||||||
| Cal MB5 is ramped up to 6kV | |||||||
| 17:50 | According to Sonny's instruction, I made "scd_clear_spars.stol" | YS | |||||
| SCI_SW_SCD_SET 0xFFF, 0x FFFF | |||||||
| SCI_SW_SCD_ENABLE_AUSTV 0xFFF, 0, 3 | |||||||
| SCI_SW_SCD_PED_UPDATE_RUN 100 | |||||||
| And I sent it. | |||||||
| >>>But, still STV in s.spi3 are still more than 30000 | |||||||
| 18:50 | send "scd_ch_coher.stol" | YS | |||||
| send "scd_ch_mask_dummy.stol" | |||||||
| send "scd_ch_mask_hot.stol" | |||||||
| send "scd_ch_mask_hot2.stol" | |||||||
| send "scd_ch_mask_hot3.stol" | |||||||
| 19:58 | send "power_on_tcd-stol" | TG | |||||
| send "hv_on_tcd_stol" | TG | ||||||
| 20:30 | Cal MB5 is dead again | TG | |||||
| 21:04 | send "hv_on_tcd_stol" | TG | |||||
| 21:21 | send "hv_on_tcd_stol" | TG | |||||
| 21:50 | Cal Mb5 Bias off | TG | |||||
| 21:52 | Cal Mb5 Hv on to 0 | TG | |||||
| 21:55 | Syncro Cal MB5 | TG | |||||
| 21:57 | Cal Mb5 Bias on | TG | |||||
| 22:02 | Cal Mb5 Hv ramping upto 750 | TG | |||||
| 22:15 | Send cal threshold of 80 | TG | |||||